# Non-Gaussianities in primordial black hole formation and induced gravitational waves

### v1

@@ III.2 Constant-roll on a quadratic potential
where $V_{0}\gg|m^{2}\varphi^{2}|$, such that the Hubble parameter can still be approximated by a constant $H\approx\sqrt{V_{0}}/\sqrt{3}M_{\mathrm{Pl}}$, while the second slow-roll parameter $\eta\equiv m^{2}/3H^{2}$ is also a constant. The equation of motion is
@@
Considering the boundary condition of $\pi(0)=\pi_{t}$, $\varphi(0)=\varphi_{t}$, the solution is
@@
which is a direct consequence of the solution (36). This is the fiducial $e$-folding number. For a perturbed universe where $\tilde{\varphi}=\varphi+\delta\varphi$, $\tilde{\pi}=\pi+\delta\pi$, $\tilde{\pi}_{t}=\pi_{t}+\delta\pi_{t}$, $\tilde{N}=N+\delta N$, we have
@@
We see that the curvature perturbation has two equivalent expressions, which is called logarithmic duality in Pi and Sasaki (2023b). Apparently it comes from the quadratic nature of the potential, which is good enough when the field excursion we study is not too long. The equivalence of the upper and lower formulas in (40) can be easily shown by the perturbed conservation law (38):
@@
In principle, the perturbation at the boundary $\delta\pi_{t}$ induced by the initial perturbation ($\delta\varphi$, $\delta\pi$) should be (38). However, in this case we do not need it after assuming that the inflaton is already the attractor solution at $\varphi_{t}$, i.e. $\pi_{t}+\lambda_{-}\varphi_{t}\approx 0$ and $\delta\pi_{t}\approx 0$. Therefore, it is more convenient to use the lower line of (40), as the second logarithm is negligible:
$\mathcal{R}\approx\frac{1}{\lambda_{-}}\ln\left(1+\frac{\delta\pi+\lambda_{+}\delta\varphi}{\pi+\lambda_{+}\varphi}\right)=\frac{1}{\lambda_{-}}\ln\left(1-\left.\lambda_{-}\frac{\delta\varphi}{\pi}\right|_{\mathrm{attractor}}\right).$ (42)
In the second step we use the fact that $\delta\varphi$ and $\varphi$ obeys the same equation of motion (33) on superhorizon scales (i.e. $k^{2}\ll H^{2}a^{2}$), when the potential is quadratic  [^6] This is a special case of the well know proportionality $\delta\varphi\propto\dot{\varphi}/H$ on superhorizon scales for any potential in single-field inflation. . Therefore, $(\delta\pi+\lambda_{+}\delta\varphi)/(\pi+\lambda_{+}\varphi)$ is a constant along each trajectory, which can be evaluated at a much later time after horizon exit when $\varphi$ is already in the attractor (but still before the boundary) Atal et al. (2019). The phase diagram is shown in Figure 2.
Slow-roll inflation and ultra-slow-roll inflation discussed in Section III.1 are special cases of (40). Slow roll is the attractor solution when $\lambda_{-}\ll 1$, $\lambda_{+}\approx 3$. Expanding (42), we get (21) with $\mathcal{R}_{g}=-\delta\varphi/\pi=H\delta\varphi/\dot{\varphi}$ and $f_{\mathrm{NL}}=-(5/3)\lambda_{-}$. In ultra-slow-roll inflation, the non-attractor solution dominates. Substituting $\lambda_{-}=0$, $\lambda_{+}=3$ into the upper line of Eq. (40), we can get (29).
@@ III.3 A general analysis on single-field inflation
We have shown how a quadratic piece of the potential (32) can give rise to the logarithmic relation of $\mathcal{R}(\delta\varphi,\delta\pi)$, Eq. (40), where $(\delta\varphi,\delta\pi)$ are evaluated on the initial spatially-flat surface after horizon crossing, which are determined by subhorizon quantum fluctuations. In general, such a potential (32) is embedded into a global potential. The initial condition of $(\varphi,\pi)$ are usually provided by a preceding (slow-roll) stage, while after the bump/plateau, the potential is usually followed by another slow-roll potential to end inflation. Before and after the junction point, the attractor solutions are usually different. This can excite the non-attractor solution transiently, which contribute to the total $\delta N$ as we discussed above. All of such contributions should be taken into account to calculate the final $\delta N$.
@@
where $\varphi_{t}$ is the junction point of the two quadratic potentials with $V(\varphi_{t})=V_{0}$, and $\varphi_{m}$ is the minimum of $V(\varphi>\varphi_{t})$. For simplicity, we set the origin to make $\varphi=0$ the maximum (minimum) of $V(\varphi<\varphi_{t})$, when $m^{2}<0$ ($m^{\prime 2}>0$). This is an extension of the Starobinsky model with a hinged linear potential Starobinsky (1992) to quadratic level, which can be further extended to a multi-exponential potential Domènech et al. (2023). Inflation ends at $\varphi_{f}$ in the second stage. A schematic figure of this piecewise potential is shown in Fig. 3. $V(\varphi)$ is continuous, but $V^{\prime}(\varphi)$ has a jump at $\varphi_{t}$. This means $\varphi^{\prime}$ is continuous, but a change in the slope means that the slow-roll solution for $\varphi<\varphi_{t}$ is not the slow-roll solution for $\varphi>\varphi_{t}$, thus $\varphi$ will decelerate/accelerate to reach the later slow-roll solution, during which the curvature perturbation will be enhanced/suppressed, and non-Gaussianity could become large.
Figure 3: Schematic pictures of the piecewise quadratic potential $V(\varphi)$ in (46), with $m_{1}^{2}<0$ (left) and $m_{2}^{2}>0$ (right). The potential is continuous at the transition point is $\varphi_{t}$, where $V(\varphi_{t})=V_{0}$, but $V^{\prime}(\varphi_{t})$ has a jump. The origin is chosen at the local maximum (upper panel) or minimum (lower panel) of $V(\varphi<\varphi_{t})$, for $m_{1}^{2}<0$ or $m_{1}^{2}>0$, respectively. At $\varphi_{t}$, the potential is continuous, while its slope may not be. The $e$-folding number is also labeled at $\varphi$, $\varphi_{t}$, and $\varphi_{f}$.
@@
where $\lambda_{\pm}$ and $\lambda^{\prime}_{\pm}$ are the characteristic roots of each piece, defined by (34). The equivalence of upper and lower expressions of each $\delta N_{i}$ are guaranteed by the perturbed conservation law (41) for each piece. Here, $\delta\varphi$ and $\delta\pi$ are the inflaton perturbation and its momentum (i.e. time derivative with respect to $-N$) on an initial spatially-flat slice around ($\varphi$, $\pi$), which should be solved from the equation of motion for $\delta\varphi$, and evaluated at a moment slightly later than the horizon exit. In slow-roll stage, typically $\delta\varphi\approx H/(2\pi)$ and $\delta\pi\approx 0$ as it decays exponentially after horizon exit, yet it is not so simple for non-attractor stages. For a concrete example of calculating $\delta\varphi$ and $\delta\pi$ in ultra-slow-roll inflation, see Jackson et al. (2023); Domènech et al. (2023). After we know $\delta\varphi$ and $\delta\pi$, the corresponding perturbations on all the later boundaries, i.e. $\delta\pi_{t}$, $\delta\pi_{e}$, etc., can be solved by the perturbed conservation law (41) for each piece, although this equation can only be solved numerically except for a few special values of $\lambda_{-}$. As we have shown before, finally all the different trajectories induced by the initial perturbation ($\delta\pi$, $\delta\varphi$) merge to the unique attractor trajectory. Since then, the $e$-folding number only depend on $\varphi$, and $\delta N$ in this stage only depend on the time difference when $\varphi$ falls on to the attractor, which is usually Gaussian. It is straightforward to extend (47) to include more segments of quadratic potentials, and for every segment of the quadratic piece, two logarithms from the early and late boundaries are present, similar to (49).
Figure 4: The schematic potential (left panel) and the phase portrait (right panel) of the “smooth transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=-0.01$, $\varphi_{m}=0$. For blue and orange trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$ (in the unit of $H$), respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Equal-$N$ lines are parallel to $\pi\propto-3\varphi$ with almost equal intervals, thus Gaussianity is well preserved.
@@
In general, the final curvature perturbation is a sum of these two terms, and the PDF of $\mathcal{R}$ is complicated when two logarithmic functions are equally important. However, there are two important limits, which can be analytically solved. When $\pi_{t}\ll\lambda_{-}^{\prime}\left(\varphi_{t}-\varphi_{m}\right)$, the velocity at the end of flat plateau is not yet the attractor of the slow-roll potential, thus a transient acceleration stage is followed. The first term of (50) dominates, so the final curvature perturbation is non-Gaussian with $f_{\mathrm{NL}}=5/2$. On the contrary, when $\pi_{t}\gg\lambda_{-}^{\prime}(\varphi_{t}-\varphi_{m})$, the velocity at the end of flat plateau surpasses the attractor velocity of the slow-roll potential, thus a transient deceleration stage is followed. The second term of (50) dominates, so the final curvature perturbation is nearly Gaussian with $f_{\mathrm{NL}}=-(6/5)\lambda_{-}^{\prime}\approx-(6/5)\eta^{\prime}$. These two limits are called “smooth transition” and “sharp transition” respectively in Cai et al. (2018), where a parameter $h\equiv 6|\lambda^{\prime}_{-}(\varphi_{t}-\varphi_{m})/\pi_{t}|$ was defined to compare the slow-roll attractor velocity $-\lambda^{\prime}_{-}(\varphi_{t}-\varphi_{m})$ and the end-of-ultra-slow-roll velocity $\pi_{t}$ at the transition point $\varphi_{t}$. To avoid confusion, we should emphasize that the comparison is between the velocities, but not between the slope of the potential.
@@
Figure 5: The schematic potential (left panel) and the phase portrait (right panel) of the “sharp transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=0.01$, $\varphi_{m}=100$ (in the unit of $H$). For the fiducial (blue) and perturbed (orange) trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$, respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Near the origin, the trajectories merge to a unique fast-roll trajectory with different $\pi$, which contribute negligibly to $\delta N$.
@@ III.4 The curvaton scenario
The curvaton scenario is a special two-field inflation model, of which the curvature perturbation is generated by the density perturbation of the curvaton field $\chi$ at its decay after inflation. This was first proposed in Moroi and Takahashi (2001); Enqvist and Sloth (2002); Lyth et al. (2003) as an alternative method of generating curvature perturbation, which was usually referred to as the outcome of the quantum fluctuation of the inflaton field $\varphi$ during inflation. In the simplest curvaton models the power spectrum is nearly scale-invariant. To generate PBHs, the power spectrum should be enhanced on a specific scale, which can be realized in the axion-curvaton model Kasuya and Kawasaki (2009); Kawasaki et al. (2013a, b); Ando et al. (2018b), non-minimal curvaton scenario Pi and Sasaki (2023a); Meng et al. (2023); Chen et al. (2023), or the multi-curvaton model Suyama and Yokoyama (2011). In all the cuvaton models, it was shown that non-Gaussianity is inevitably enhanced compared to the slow-roll inflation, and the non-linear parameter is Bartolo et al. (2004); Enqvist and Nurmi (2005); Sasaki et al. (2006)  [^7] For simplicity, we only consider a quadratic potential for the curvaton $\chi$, and the decay is supposed to be instantaneous. For more details of these assumptions, see Sasaki et al. (2006).
@@
where $r$ is approximately the cosmological parameter of the curvaton $\chi$ at its decay. Apparently such a nominal $f_{\mathrm{NL}}$ is not enough to calculate the PBH abundance, as $f_{\mathrm{NL}}$ is always larger than 1.
We will skip the model building and the dynamics of the curvaton field, but focus on its fully non-Gaussian curvature perturbation on the uniform-density slice, which is derived in Sasaki et al. (2006) for a Gaussian curvaton field perturbation in the sudden decay approximation:
@@
where dots are higher order terms of $\mathcal{O}(r^{3})$. We see that $\zeta$ has the familiar form of the quadratic local non-Gaussianity (29), with $\zeta_{g}=(2/3)r\delta\chi/\chi$ and $f_{\text{NL}}=5/(4r)$. Cubic and higher terms are suppressed in the $r\to 0$ limit, which makes it possible to truncate (56) at quadratic level when $f_{\mathrm{NL}}\gg 1$ (i.e. $r\ll 1$). Therefore, counterintuitively, such a quadratic expression (56) holds for large $f_{\mathrm{NL}}$, towards the limit of a $\chi^{2}$-distribution.
The opposite limit $r\to 1$ is more interesting and attracts much attention recently. In this limit, we simply have $P=2(Ar)^{2}$, $K=(Ar)^{2/3}/2$, and
$\zeta=\frac{2}{3}\ln A=\frac{2}{3}\ln\left(1+\frac{\delta\chi}{\chi}\right).$ (57)
@@
This means that the curvaton scenario can suppress the formation of PBHs in the $r\to 1$ limit, which is a natural model to avoid the PBH overproduction problem when interpreting the nHz stochastic GWs as induced GWs Franciolini et al. (2023a).

### v2

@@ III.2 Constant-roll on a quadratic potential
shown in Fig.2. We assume $V_{0}\gg|m^{2}\varphi^{2}|$, such that the Hubble parameter can still be approximated by a constant $H\approx\sqrt{V_{0}}/\sqrt{3}M_{\mathrm{Pl}}$, while the second slow-roll parameter $\eta\equiv m^{2}/3H^{2}$ is also a constant. The equation of motion is
@@
Considering the boundary conditions $\pi(0)=\pi_{t}$, $\varphi(0)=\varphi_{t}$, the special solution is
@@
where both sides equal to $e^{\lambda_{+}\lambda_{-}N}=e^{3\eta N}$, a direct consequence of the solution (36). For a perturbed universe where $\tilde{\varphi}=\varphi+\delta\varphi$, $\tilde{\pi}=\pi+\delta\pi$, $\tilde{\pi}_{t}=\pi_{t}+\delta\pi_{t}$, $\tilde{N}=N+\delta N$, we have
@@
We see that the curvature perturbation has two equivalent expressions, which is called logarithmic duality in Ref. Pi and Sasaki (2023b). Apparently it comes from the quadratic nature of the potential, which is a good approximation when the field excursion we study is not too long. The equivalence of the upper and lower formulas in (40) can be easily shown by the perturbed version of conservation law (38):
@@
In principle, the perturbation at the boundary $\delta\pi_{t}$ induced by the initial perturbation ($\delta\varphi$, $\delta\pi$) should be solved by (41). However, in this case we do not need it after assuming that the inflaton is already the attractor solution at $\varphi_{t}$, i.e. $\pi_{t}+\lambda_{-}\varphi_{t}\approx 0$ and $\delta\pi_{t}\approx 0$. Therefore, it is more convenient to use the lower line of (40), as the second logarithm is negligible:
$\mathcal{R}\approx\frac{1}{\lambda_{-}}\ln\left(1+\frac{\delta\pi+\lambda_{+}\delta\varphi}{\pi+\lambda_{+}\varphi}\right)=\frac{1}{\lambda_{-}}\ln\left(1-\left.\lambda_{-}\frac{(\delta\varphi)_{k}}{\pi}\right|_{a\gg k/H}\right).$ (42)
In the second step we use the fact that $\delta\varphi$ and $\varphi$ obeys the same equation of motion (33) on superhorizon scales (i.e. $k^{2}\ll H^{2}a^{2}$), when the potential is quadratic  [^6] This is a special case of the well know proportionality $\delta\varphi\propto\dot{\varphi}/H$ on superhorizon scales for any potential in single-field inflation. . Therefore, $(\delta\pi+\lambda_{+}\delta\varphi)/(\pi+\lambda_{+}\varphi)$ is a constant along each trajectory, which can be evaluated at a much later time after horizon exit (marked by $a\gg k/H$) when $\varphi$ is already in the attractor (but still before the boundary) Atal et al. (2019). Note that $(\delta\varphi)_{k}/\pi$ in (42) was stretched out of the horizon at $a_{k}=k/H$, which is now on a much larger scale of $a/k\gg 1/H$. It can be calculated by evolving the solution of Mukhanov-Sasaki equation to late time $a\gg k/H$, which should not be confused with the “new” perturbations exiting the horizon at that moment. The phase diagram is shown in Fig.2.
Slow-roll inflation and ultra-slow-roll inflation discussed in Section III.1 are special cases of (40). Slow roll is the attractor solution when $\lambda_{-}\ll 1$, $\lambda_{+}\approx 3$. Expanding (42), we get (21) with $\mathcal{R}_{g}=-\delta\varphi/\pi=H\delta\varphi/\dot{\varphi}$ and $f_{\mathrm{NL}}=-(5/6)\lambda_{-}$. In ultra-slow-roll inflation, the non-attractor solution dominates. Substituting $\lambda_{-}=0$, $\lambda_{+}=3$ into the upper line of Eq. (40), we can get (29).
@@ III.3 A general analysis on single-field inflation
We have shown how a quadratic piece of the potential (32) can give rise to the logarithmic relation (40) of $\mathcal{R}(\delta\varphi,\delta\pi)$. The Gaussian variables $(\delta\varphi,\delta\pi)$ originate from quantum fluctuations on subhorizon scales, which should be evaluated on the initial spatially-flat surface after the horizon exit. In general, such a potential (32) is a segment of a global potential. In a typical three-segment model, the initial condition of $(\varphi,\pi)$ are usually provided by a preceding (slow-roll) stage, while after the bump/plateau, the potential is usually followed by another slow-roll potential to end inflation. Before and after the junction point, the attractor solutions are usually different. This can excite the non-attractor solution transiently, which contribute to the total $\delta N$ as we discussed above. All of such contributions should be taken into account to calculate the final $\delta N$.
@@
where $\varphi_{t}$ is the junction point of the two quadratic potentials with $V(\varphi_{t})=V_{0}$, and $\varphi_{m}$ is the minimum of $V(\varphi>\varphi_{t})$. For simplicity, we set the origin to make $\varphi=0$ the maximum (minimum) of $V(\varphi<\varphi_{t})$, when $m^{2}<0$ ($m^{2}>0$). This is an extension of the Starobinsky model with a hinged linear potential Starobinsky (1992) to quadratic level, which can be further extended to a multi-exponential potential Domènech et al. (2023). Inflation ends at $\varphi_{e}$ in the second stage. A schematic figure of this piecewise potential is shown in Fig. 3. $V(\varphi)$ is continuous, but $V^{\prime}(\varphi)$ has a jump at $\varphi_{t}$. This means $\varphi^{\prime}$ is continuous, but a change in the slope means that the attractor solution for $\varphi<\varphi_{t}$ is not the attractor solution for $\varphi>\varphi_{t}$, thus $\varphi$ will decelerate/accelerate to reach the later attractor solution, during which the curvature perturbation will be enhanced/suppressed, and large non-Gaussianity could be generated.
Figure 3: Schematic pictures of the piecewise quadratic potential $V(\varphi)$ in (43), with $m^{2}<0$ (left) and $m^{2}>0$ (right). The potential is continuous at the transition point $\varphi_{t}$, where $V(\varphi_{t})=V_{0}$, but $V^{\prime}(\varphi_{t})$ has a jump. The origin is chosen at the local maximum (left panel) or minimum (right panel) of $V(\varphi<\varphi_{t})$, for $m^{2}<0$ or $m^{2}>0$, respectively. The $e$-folding numbers are labeled at $\varphi$, $\varphi_{t}$, and $\varphi_{e}$.
@@
where $\lambda_{\pm}$ and $\lambda^{\prime}_{\pm}$ are the characteristic roots of each piece, defined by (34). The equivalence of upper and lower expressions of each $\delta N_{i}$ are guaranteed by the perturbed conservation law (41) for each piece. Here, $\delta\varphi$ and $\delta\pi$ are the inflaton perturbation and its momentum (i.e. time derivative with respect to $-N$) on an initial spatially-flat slice around ($\varphi$, $\pi$), which should be solved from the equation of motion for $\delta\varphi$, and evaluated at a moment slightly later than the horizon exit. In slow-roll stage, typically $\delta\varphi\approx H/(2\pi)$ and $\delta\pi\approx 0$ as $\delta\pi$ decays exponentially after horizon exit, yet it is not so simple for non-attractor stages. For a concrete example of calculating $\delta\varphi$ and $\delta\pi$ in ultra-slow-roll inflation, see Jackson et al. (2023); Domènech et al. (2023). After we know $\delta\varphi$ and $\delta\pi$, the corresponding perturbations on all the later boundaries, i.e. $\delta\pi_{t}$, $\delta\pi_{e}$, etc., can be solved by the perturbed conservation law (41) for each piece, although this equation can only be solved numerically except for a few special values of $\lambda_{-}$ listed in Ref. Pi and Sasaki (2023b). As we have shown before, finally all the different trajectories induced by the initial perturbation ($\delta\pi$, $\delta\varphi$) merge to the unique attractor trajectory. Since then, the $e$-folding number only depends on $\varphi$, and $\delta N$ in this stage only depends on the time difference when $\varphi$ falls on to the attractor, which is usually Gaussian. It is straightforward to extend (44) to include more segments of quadratic potentials, and for every segment of the quadratic piece, two logarithms from the early and late boundaries are present, similar to (46).
Figure 4: The schematic potential (left panel) and the phase portrait (right panel) of the “smooth transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=-0.01$, $\varphi_{m}=0$. For fiducial (blue) and perturbed (orange) trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$ (in the unit of $H$), respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Equal-$N$ lines are parallel to the fast-roll streams (i.e. $\pi\propto-3\varphi$) with almost equal intervals, which are not drawn for visual clearance. This clearly shows that the contribution to $\delta N$ from the ultra-slow-roll stage is negligible, and the contribution from the later stage preserved slow-roll-type Gaussianity.
@@
In general, the final curvature perturbation is a sum of these two terms, and the PDF of $\mathcal{R}$ can only be solved numerically when two logarithmic functions are equally important. However, it can be solved analytically when one of the logarithms dominates. When $\pi_{t}\gg\lambda_{-}^{\prime}(\varphi_{t}-\varphi_{m})$, the velocity at the end of flat plateau surpasses the attractor velocity of the slow-roll potential, thus a transient deceleration stage is followed. The second term of (47) dominates, so the final curvature perturbation is nearly Gaussian with $f_{\mathrm{NL}}=-(6/5)\lambda_{-}^{\prime}\approx-(6/5)\eta^{\prime}$. On the contrary, when $\pi_{t}\ll\lambda_{-}^{\prime}\left(\varphi_{t}-\varphi_{m}\right)$, the velocity at the end of flat plateau is not yet the attractor of the slow-roll potential, thus a transient acceleration stage is followed. The first term of (47) dominates, so the final curvature perturbation is non-Gaussian with $f_{\mathrm{NL}}=5/2$. These two limits are called “smooth transition” and “sharp transition” respectively in Cai et al. (2018), where a parameter $h\equiv 6|\lambda^{\prime}_{-}(\varphi_{t}-\varphi_{m})/\pi_{t}|$ was defined to compare the slow-roll attractor velocity $-\lambda^{\prime}_{-}(\varphi_{t}-\varphi_{m})$ and the realistic end-of-ultra-slow-roll velocity $\pi_{t}$ at the transition point $\varphi_{t}$. To avoid confusion, we should emphasize that the comparison is between the velocities, but not between the slopes of the potentials.
@@
Figure 5: The schematic potential (left panel) and the phase portrait (right panel) of the “sharp transition” from ultra-slow-roll to slow-roll, with $\eta^{\prime}=0.01$, $\varphi_{m}=100$ (in the unit of $H$). For the fiducial (blue) and perturbed (orange) trajectories we set $\pi=9.15,~{}9.05$ at $\varphi=-3$, respectively. To show the difference of the trajectories clearly, $\delta\pi$ is exaggerated. Near the origin, the trajectories merge to a unique fast-roll trajectory with different $\pi$, which contribute negligibly to $\delta N$ in the subsequent evolution.
@@ III.4 The curvaton scenario
The curvaton scenario is a special two-field inflation model, of which the curvature perturbation is generated by the density perturbation of the curvaton field $\chi$ at its decay after inflation. This was first proposed in Moroi and Takahashi (2001); Enqvist and Sloth (2002); Lyth et al. (2003) as an alternative method of generating curvature perturbation, which was usually referred to as the outcome of the quantum fluctuation of the inflaton field $\varphi$ during inflation. In the simplest curvaton models the power spectrum is nearly scale-invariant. To generate PBHs, the power spectrum should be enhanced on a specific scale, which can be realized in the axion-curvaton model Kasuya and Kawasaki (2009); Kawasaki et al. (2013a, b); Ando et al. (2018b), non-minimal curvaton scenario Pi and Sasaki (2023a); Meng et al. (2023); Chen et al. (2023), or the multi-curvaton model Suyama and Yokoyama (2011). In the cuvaton scenario, it was shown that non-Gaussianity is inevitably enhanced compared to the slow-roll inflation, and the non-linear parameter is Bartolo et al. (2004); Enqvist and Nurmi (2005); Sasaki et al. (2006) [^7] For simplicity, we only consider a quadratic potential for the curvaton $\chi$, and the decay is supposed to be instantaneous. For more details of these assumptions, see Ref. Sasaki et al. (2006).
@@
where $r$ is roughly the cosmological parameter of the curvaton $\chi$ at its decay. Apparently such a nominal $f_{\mathrm{NL}}$ is not enough to calculate the PBH abundance, as $f_{\mathrm{NL}}$ is always larger than 1.
We will skip the model building and the dynamics of the curvaton field, but focus on its fully non-Gaussian curvature perturbation on the uniform-density slice, which is derived in Ref. Sasaki et al. (2006) for a Gaussian curvaton field perturbation in the sudden decay approximation:
@@
where dots are higher order terms of $\mathcal{O}(r^{3})$. We see that $\zeta$ has the familiar form of the quadratic local non-Gaussianity (29), with $\zeta_{g}=(2/3)r\delta\chi/\chi$ and $f_{\text{NL}}=5/(4r)$. Cubic and higher-order terms are suppressed when $r\ll 1$, which makes it possible to truncate (53) at quadratic level even when $f_{\mathrm{NL}}\gg 1$. Therefore, counterintuitively, such a quadratic expression (53) holds for large $f_{\mathrm{NL}}$, towards the limit of a $\chi^{2}$-distribution.
The opposite limit $r\to 1$ is more complicated and attracts much attention recently. In this limit, we simply have $P=2(Ar)^{2}$, $K=(Ar/2)^{2/3}$, and
$\zeta=\frac{1}{3}\ln A=\frac{2}{3}\ln\left(1+\frac{\delta\chi}{\chi}\right).$ (54)
@@
This means that the curvaton scenario can suppress the formation of PBHs in the $r\to 1$ limit, which is a natural model to avoid the overproduction of PBHs when interpreting the nHz stochastic GWs as induced GWs Franciolini et al. (2023a).